OpenStudy (anonymous):

simplify of the expression (b5)4.

OpenStudy (lgbasallote):

multiply your exponents

OpenStudy (anonymous):

they are powers.. not normal numbers by it

OpenStudy (anonymous):

so it would be (B20)

OpenStudy (lgbasallote):

B?

OpenStudy (anonymous):

no the numbers are like small and above.. not like a number by B

OpenStudy (anonymous):

power of a power rule: multiply the exponents so your right: b^20

OpenStudy (lgbasallote):

i know... \[\large (a^m)^n = a^{mn}\] myltiply your exponents like i said =)))

OpenStudy (anonymous):

okay :D thanks... what about his problem

OpenStudy (anonymous):

c2 • c9

OpenStudy (anonymous):

same thing? multiply them?

OpenStudy (lgbasallote):

\[\large a^m \times a^n = a^{(m+n)}\]

OpenStudy (anonymous):

alrigh tthank you!

OpenStudy (anonymous):

what about (d3)5 ?

OpenStudy (anonymous):

power of a power, so multiply the xponents

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

denote the exponents with ^ because it's confusing.. i.e. x^2 is x squared

OpenStudy (anonymous):

alright il remember that next time